# On the Natural Robustness of Vision-Language Models Against Visual Perception Attacks in Autonomous Driving

### v2

@@ 4.1 Experimental Setup
Although the fine-tuned models were trained on AV-specific data, the AEs remained unseen during training, allowing a reliable evaluation of each V2LM’s resilience to attacks on TSR, ALC, and VD. To achieve this, three types of black-box attacks were implemented. The first type of attack involves adversarial manipulation of traffic signs. In the study by Zhong et al. [72], shadows are utilized to conduct attacks on TSR algorithms, as shown in Figure 4.
@@
Table 2: Comparison of V2LM Performance in AV Tasks. Green Color shows Improvement of Tandem over Solo Model. LLaVA-13B is abbreviated for LLaVA-13B-LoRA.
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 95.93% 97.38% 1.45% 94.62% 91.22% -3.40% 94.94% 91.08% -3.86% 94.31% 91.37% -2.94% LLaVA-13B 97.15% 98.14% 0.99% 96.09% 93.21% -2.88% 96.25% 92.35% -3.90% 95.94% 94.10% -1.84% MoE-LLaVA 95.24% 96.91% 1.67% 94.73% 95.78% 1.05% 94.51% 95.26% 0.75% 94.96% 96.31% 1.35% MobileVLM 88.19% 90.34% 2.15% 87.68% 90.18% 2.50% 86.17% 90.73% 4.56% 89.25% 89.65% 0.40% Qwen-7B 95.80% 94.13% -1.67% 95.78% 94.27% -1.51% 95.96% 94.63% -1.33% 95.64% 93.88% -1.76% NVILA-8B 99.04% 99.14% 0.1% 99.10% 99.20% 0.1% 99.11% 99.19% 0.08% 99.13% 99.14% 0.01% ALC LLaVA-7B 98.11% 95.41% -2.70% 98.05% 95.24% -2.81% 98.20% 95.21% -2.99% 97.91% 95.29% -2.62% LLaVA-13B 99.30% 98.31% -0.99% 99.46% 97.85% -1.61% 99.46% 98.32% -1.14% 99.47% 97.40% -2.07% MoE-LLaVA 96.86% 92.53% -4.33% 96.23% 91.83% -4.40% 97.63% 92.37% -5.26% 96.85% 91.30% -5.55% MobileVLM 84.41% 86.61% 2.20% 87.68% 89.53% 1.85% 86.17% 88.93% 2.76% 89.26% 90.10% 0.84% Qwen-7B 93.75% 93.91% 0.16% 93.83% 93.76% -0.07% 94.39% 94.48% 0.09% 93.27% 93.05% -0.22% NVILA-8B 99.51% 99.41% -0.1% 99.45% 99.52% 0.07% 99.34% 99.44% 0.1% 99.55% 99.61% 0.06% VD LLaVA-7B 95.84% 91.90% -3.94% 94.56% 92.82% -1.74% 95.12% 94.71% -0.41% 94.01% 91.21% -2.80% LLaVA-13B 97.05% 98.96% 1.91% 96.76% 95.97% -0.79% 97.33% 97.88% 0.55% 96.21% 94.14% -2.07% MoE-LLaVA 95.52% 93.42% -2.10% 94.78% 92.30% -2.48% 95.31% 93.83% -1.48% 94.27% 90.82% -3.45% MobileVLM 86.61% 88.01% 1.40% 88.42% 89.38% 0.96% 86.76% 87.52% 0.76% 90.15% 91.34% 1.19% Qwen-7B 96.02% 95.01% -1.01% 96.55% 96.17% -0.38% 97.76% 97.35% -0.41% 94.91% 95.01% 0.10% NVILA-8B 99.93% 99.95% 0.02% 99.83% 99.84% 0.01% 99.81% 99.79% -0.02% 99.88% 99.91% 0.03%
Perception GuardPerception SystemV2LMTSRALCVDDynamic ObjectTraffic SignLane DetectionDetectionTrackingPredictionDetectionRecognitionDetectionTrackingIroadviewOutput+
@@
The second type of attack targets the ALC mechanism of AVs. In the study by Sato et al. [51], the Dirty Road Patch (DRP) attack framework specifically targets ALC systems in AVs, exploiting vulnerabilities in deep learning-based lane detection. This method employs an optimization-based approach to systematically generate these patches, as illustrated in Figure 4, considering real-world conditions such as lighting and camera angles to ensure effectiveness across different environmental scenarios. The optimized DRPs cause the AV to make incorrect steering decisions, which were demonstrated to be highly successful in real-world driving scenarios with a success rate exceeding 97.5%
@@
Figure 4 shows the third type of attack which focuses on the adversarial camouflage of vehicles. Zhou et al. [73] propose a physical adversarial attack known as the Robust and Accurate UV-map-based Camouflage Attack (RAUCA) to deceive VD algorithms such as YOLOv3 [48]. It employs a technique utilizing a differentiable neural renderer, which allows for the optimization of adversarial camouflages through gradient back-propagation, enhancing both the robustness and precision of the attacks under varying environmental conditions. Their method achieved an attack success rate of 97.48% on the target detection models, demonstrating the significant vulnerability of these systems to such sophisticated camouflage attacks.
(a) Shadow Attack
(b) DRP Attack
(c) RAUCA
Figure 4: Examples of adversarial attacks targeting AV perception.
4.2 RQ1: Fine-Tuning Increases Detection Performance
Table 1 shows VLMs’ zero-shot performance on the test dataset, which performed poorly in ALC and TSR tasks, indicating difficulties in these specific AV applications. Although the models demonstrated decent performance in the VD task, this success may partly be due to their pre-training on large, diverse image datasets, which likely enhanced their general visual recognition capabilities. However, they struggled to detect objects in images taken during darkness, rain, or when the objects were far away, showing their difficulty in maintaining high performance under varying and challenging environmental conditions despite their general visual understanding capabilities.
@@ 4.2 RQ1: Fine-Tuning Increases Detection Performance
LLaVA -7B LLaVA -13B MoE -LLaVA Mobile VLM Qwen -7B NVILA -8B 000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.81111IoU ValuePre-trainedFine-Tuned (Solo)Fine-Tuned (Tandem)
@@
To improve their performance, these models were fine-tuned using the same training datasets and prompts as before, and then their performance was re-evaluated with the same test dataset. Table 2 shows notable accuracy improvements for all tasks after fine-tuning. For ALC, accuracy increased from 20.71%–50.91% to 86.61%–99.51%. For TSR, accuracy saw a substantial rise from 1.19%–31.24% to 90.34%–99.14%. In the VD task, accuracy improved from 60.64%–92.06% to 88.01
@@
These findings suggest that although they initially underperformed on AV tasks, fine-tuning them with relevant datasets can lead to substantial performance improvements, showing their potential utility in AV applications. Intersection over Union (IoU) measures the overlap between the predicted output and the ground truth in tasks like object detection. A higher IoU indicates more accurate localization, making it a crucial metric for evaluating model performance in AV perception tasks. After fine-tuning, the IoU values improved significantly, with LLaVA-7B increasing from 0.041 to 0.964, LLaVA-13B-LoRA from 0.063 to 0.987, MoE-LLaVA from 0.038 to 0.946, MobileVLM from 0.013 to 0.926, Qwen-7B from 0.018 to 0.9578, and NVILA-8B from 0.183 to 0.9768, as shown in Figure 5.
4.3 RQ2: V2LMs Demonstrate Robustness under Attacks
Table 3: Performance of DNN models evaluated on benign (B) and unseen adversarial (A) datasets across AV tasks: TSR, ALC, and VD. The difference (Diff) indicates the accuracy degradation caused by adversarial attacks.
Task Model Type Accuracy F1-Score Precision Recall
TSR Yolov5-cls A 55.82% 58.10% 60.56% 55.82%
B 94.88% 93.44% 94.88% 93.92%
Diff -39.06% -35.34% -34.32% -38.10%
ALC CLRerNet A 50.51% 44.86% 47.66% 50.51%
B 90.91% 90.83% 91.44% 90.91%
Diff -40.40% -45.97% -43.78% -40.40%
VD Yolov5-dt A 62.27% 57.44% 51.37% 64.11%
B 97.01% 96.60% 96.66% 97.01%
Diff -34.74% -39.16% -45.29% -32.91%
@@ 4.3 RQ2: V2LMs Demonstrate Robustness under Attacks
Adversarial attacks pose a serious threat to deep learning-based AV perception systems. Prior works have shown that DNN models suffer significant performance degradation across core AV tasks, including TSR, ALC, and VD, as summarized in Table 5. For example, the GTSRB-CNN model accuracy collapses to just 1.77% under the Shadow Attack [72], and even adversarial training—one of the most prominent defense methods—only modestly improves accuracy to 25.57%. Similarly, for ALC, OpenPilot-ALC performance degrades dramatically to 2.50% under the DRP Attack [51], with established defense strategies such as JPEG compression [20], Gaussian noise addition [64], and autoencoder-based denoising [19] achieving negligible improvements (around 3%). In VD, YOLOv3 experiences a drastic accuracy drop to 2.52% under RAUCA Attack [73], with no effective defense method proposed.
@@
Building on these findings, we conduct our own robustness evaluation. To provide a fair and rigorous comparison, we evaluated traditional task-specific DNNs—YOLOv5-cls [26, 47] (TSR), CLRerNet [21] (ALC), and YOLOv5-dt (VD)—on benign and previously unseen adversarial datasets (Table 3). Our results confirm severe performance degradation, with accuracy reductions of 39.06% for TSR, 40.40% for ALC, and 32.91% for VD, clearly highlighting their vulnerability to novel attacks.
In contrast, our evaluations of V2LMs under the same unseen adversarial conditions reveal inherently superior robustness. Specifically, V2LMs experienced substantially smaller accuracy drops: only 8.62%–15.81% for TSR, 4.94%–14.32% for ALC, and 5.23%–7.06% for VD, consistently maintaining high adversarial accuracy without additional defenses. These results strongly support the potential of V2LMs to enhance robustness and reliability of AV perception systems.
Table 4: Comparison of V2LM Performance against AEs. Green Color shows Improvement of Tandem over Solo Model. LLaVA-13B is abbreviated for LLaVA-13B-LoRA.
Accuracy F1-Score Precision Recall Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 80.12% 86.51% 6.39% 86.44% 85.89% -0.55% 85.24% 86.12% 0.88% 87.69% 85.67% -2.02% LLaVA-13B 86.53% 89.01% 2.48% 86.08% 88.95% 2.87% 85.13% 89.31% 4.18% 87.06% 88.60% 1.54% MoE-LLaVA 80.03% 81.52% 1.49% 80.75% 82.34% 1.59% 82.13% 83.26% 1.13% 79.43% 81.45% 2.02% MobileVLM 79.57% 79.13% -0.44% 77.55% 79.88% 2.33% 78.02% 80.75% 2.73% 77.10% 79.04% 1.94% Qwen-7B 82.83% 81.66% -1.17% 83.16% 83.43% 0.27% 85.73% 86.52% 0.79% 80.57% 80.77% 0.20% NVILA-8B 86.04% 91.12% 5.08% 87.15% 91.10% 3.95% 90.28% 91.15% 0.87% 86.09% 91.10% 5.01% ALC LLaVA-7B 86.07% 84.83% -1.24% 87.01% 85.38% -1.63% 86.68% 84.65% -2.03% 87.53% 86.13% -1.40% LLaVA-13B 90.05% 86.69% -3.36% 90.69% 87.38% -3.31% 89.67% 86.34% -3.33% 91.75% 88.46% -3.29% MoE-LLaVA 82.54% 82.26% -0.28% 83.27% 83.06% -0.21% 84.46% 82.49% -1.97% 82.13% 83.65% 1.52% MobileVLM 78.21% 78.23% 0.02% 80.92% 81.21% 0.29% 83.16% 81.96% -1.20% 78.80% 80.49% 1.69% Qwen-7B 88.81% 86.57% -2.24% 89.48% 85.48% -4% 91.06% 89.61% -1.45% 88.01% 86.57% -1.44% NVILA-8B 99.25% 99.14% –0.11% 99.13% 99.10% –0.03% 99.01% 99.11% +0.10% 99.00% 99.10% +0.10% VD LLaVA-7B 89.23% 88.02% -1.21% 87.84% 86.59% -1.25% 89.06% 87.89% -1.17% 86.67% 85.33% -1.34% LLaVA-13B 91.82% 90.09% -1.73% 90.72% 90.41% -0.31% 90.10% 91.03% 0.93% 91.35% 89.80% -1.55% MoE-LLaVA 88.46% 86.54% -1.92% 87.16% 86.21% -0.95% 88.28% 86.77% -1.51% 86.07% 85.67% -0.40% MobileVLM 80.13% 80.15% 0.02% 82.88% 82.97% 0.09% 84.14% 83.81% -0.33% 81.67% 82.16% 0.49% Qwen-7B 89.05% 89.06% 0.01% 90.15% 90.57% 0.42% 92.83% 92.77% -0.06% 87.60% 88.46% 0.86% NVILA-8B 99.81% 99.83% 0.02% 99.82% 99.83% 0.01% 99.87% 99.81% -0.06% 99.81% 99.87% 0.06%
4.4 RQ3: Tandem V2LMs Provide Similar Performance at Lower Memory Footprint
This section evaluates whether a single tandem V2LM can handle multiple defense tasks in detecting AEs across AV systems, compared to using separate solo V2LMs for each task. To assess this, VLMs are fine-tuned for three tasks — TSR, ALC, and VD — simultaneously using the same prompts. After fine-tuning, both solo and tandem V2LMs were then evaluated on each task to determine its performance on the same test data.
Table 2 presents the evaluation results for the tandem design in non-adversarial scenarios, where a single V2LM was trained to handle all three tasks simultaneously. The results reveal that the tandem design achieves high accuracy across perception tasks, often matching or surpassing the performance of solo models. This shows the effectiveness of the tandem design in maintaining robust performance across diverse tasks. Table 4 further demonstrates the resilience of V2LMs in AV tasks under adversarial conditions.
@@ 4.4 RQ3: Tandem V2LMs Provide Similar Performance at Lower Memory Footprint
MobileVLM, MoE-LLaVA, LLaVA-7B, NVILA-8B, Qwen-VL, and LLaVA-13B-LoRA allocated 6.03GB, 11.21GB, 13.56GB, 15.23GB, 16.58GB, and 26.15GB of storage, respectively. These results suggest that a single tandem V2LM can generalize well across multiple tasks, providing robust performance comparable to the solo design, which requires 3x more storage for separate models. The tandem design offers a significant advantage in efficiency by maintaining similar performance while requiring much less storage.
Table 5: Traditional Models Accuracy under Adversarial Attacks and Defense Strategies.
@@
ALC OpenPilot- ALC DRP [51] 2.50% JPEG Compression [20] Bit-Depth Reduction [65] Gaussian Noise [64] Median Blurring [65] Autoencoder [19] $\approx$3(no effective improvement)

### v3

@@ IV-A Experimental Setup
Although the fine-tuned models were trained on AV-specific data, the AEs remained unseen during training, allowing a reliable evaluation of each V2LM’s resilience to attacks on TSR, ALC, and VD. To achieve this, three types of black-box attacks were implemented. The first type of attack involves adversarial manipulation of traffic signs. In the study by Zhong et al. [56], shadows are utilized to conduct attacks on TSR algorithms, as shown in figure 2a.
@@
TABLE III: Performance of DNN models on benign (B) and unseen adversarial (A). Diff is the accuracy/F1 drop due to attacks.
Task Model Metric A B Diff
TSR YOLOv5-cls Acc. 55.82% 94.88% -39.06%
F1-Score 58.10% 93.44% -35.34%
ViT-small Acc. 23.12% 96.20% -73.08%
F1-Score 23.12% 96.36% -73.24%
ALC CLRerNet Acc. 50.51% 90.91% -40.40%
F1-Score 44.86% 90.83% -45.97%
VD YOLOv5-dt Acc. 62.27% 97.01% -34.74%
F1-Score 57.44% 96.60% -39.16%
@@
The second type of attack targets the ALC mechanism of AVs. In the study by Sato et al. [37], the Dirty Road Patch (DRP) attack framework specifically targets ALC systems in AVs, exploiting vulnerabilities in deep learning-based lane detection. This method employs an optimization-based approach to systematically generate these patches, as illustrated in figure 2b, considering real-world conditions such as lighting and camera angles to ensure effectiveness across different environmental scenarios. The optimized DRPs cause the AV to make incorrect steering decisions, which were demonstrated to be highly successful in real-world driving scenarios with a success rate exceeding 97.5%
@@
Figure 2c shows the third type of attack which focuses on the adversarial camouflage of vehicles. Zhou et al. [57] propose a physical adversarial attack known as the Robust and Accurate UV-map-based Camouflage Attack (RAUCA) to deceive VD algorithms such as YOLOv3 [34]. It employs a technique utilizing a differentiable neural renderer, which allows for the optimization of adversarial camouflages through gradient back-propagation, enhancing both the robustness and precision of the attacks under varying environmental conditions. Their method achieved an attack success rate of 97.48% on the target detection models, demonstrating the significant vulnerability of these systems to such sophisticated camouflage attacks.
TABLE IV: comparison of V2LM performance: Benign vs. AEs. Green shows improvement of Tandem over Solo.
Benign AEs Accuracy F1-Score Accuracy F1-Score Task Model Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff Solo Tandem Diff TSR LLaVA-7B 95.93% 97.38% 1.45% 94.62% 91.22% -3.40% 80.12% 86.51% 6.39% 86.44% 85.89% -0.55% LLaVA-13B 97.15% 98.14% 0.99% 96.09% 93.21% -2.88% 86.53% 89.01% 2.48% 86.08% 88.95% 2.87% MoE-LLaVA 95.24% 96.91% 1.67% 94.73% 95.78% 1.05% 80.03% 81.52% 1.49% 80.75% 82.34% 1.59% MobileVLM 88.19% 90.34% 2.15% 87.68% 90.18% 2.50% 79.57% 79.13% -0.44% 77.55% 79.88% 2.33% Qwen-VL 95.80% 94.13% -1.67% 95.78% 94.27% -1.51% 82.83% 81.66% -1.17% 83.16% 83.43% 0.27% NVILA-8B 99.04% 99.14% 0.1% 99.10% 99.20% 0.1% 86.04% 91.12% 5.08% 87.15% 91.10% 3.95% ALC LLaVA-7B 98.11% 95.41% -2.70% 98.05% 95.24% -2.81% 86.07% 84.83% -1.24% 87.01% 85.38% -1.63% LLaVA-13B 99.30% 98.31% -0.99% 99.46% 97.85% -1.61% 90.05% 86.69% -3.36% 90.69% 87.38% -3.31% MoE-LLaVA 96.86% 92.53% -4.33% 96.23% 91.83% -4.40% 82.54% 82.26% -0.28% 83.27% 83.06% -0.21% MobileVLM 84.41% 86.61% 2.20% 87.68% 89.53% 1.85% 78.21% 78.23% 0.02% 80.92% 81.21% 0.29% Qwen-VL 93.75% 93.91% 0.16% 93.83% 93.76% -0.07% 88.81% 86.57% -2.24% 89.48% 85.48% -4% NVILA-8B 99.51% 99.41% -0.1% 99.45% 99.52% 0.07% 99.25% 99.14% -0.11% 99.13% 99.10% -0.03% VD LLaVA-7B 95.84% 91.90% -3.94% 94.56% 92.82% -1.74% 89.23% 88.02% -1.21% 87.84% 86.59% -1.25% LLaVA-13B 97.05% 98.96% 1.91% 96.76% 95.97% -0.79% 91.82% 90.09% -1.73% 90.72% 90.41% -0.31% MoE-LLaVA 95.52% 93.42% -2.10% 94.78% 92.30% -2.48% 88.46% 86.54% -1.92% 87.16% 86.21% -0.95% MobileVLM 86.61% 88.01% 1.40% 88.42% 89.38% 0.96% 80.13% 80.15% 0.02% 82.88% 82.97% 0.09% Qwen-VL 96.02% 95.01% -1.01% 96.55% 96.17% -0.38% 89.05% 89.06% 0.01% 90.15% 90.57% 0.42% NVILA-8B 99.93% 99.95% 0.02% 99.83% 99.84% 0.01% 99.81% 99.83% 0.02% 99.82% 99.83% 0.01%
@@
LLaVA -7B LLaVA -13B MoE -LLaVA Mobile VLM Qwen -7B NVILA -8B 00.20.20.40.40.60.60.80.811IoU ValuePre-trainedFine-Tuned (Solo)Fine-Tuned (Tandem)
@@
IV-B RQ1: Fine-Tuning Increases Detection Performance
Table II shows VLMs’ zero-shot performance on the test dataset, which performed poorly in ALC and TSR tasks, indicating difficulties in these specific AV applications. Although the models demonstrated decent performance in the VD task, this success may partly be due to their pre-training on large, diverse image datasets, which likely enhanced their general visual recognition capabilities. However, in the VD task, they struggled to detect vehicles in images taken at night, in the rain, or at long distances (highlighting their limitations under challenging environmental conditions despite strong general visual recognition).
@@ IV-B RQ1: Fine-Tuning Increases Detection Performance
To improve their performance, these models were fine-tuned using the same training datasets and prompts as before, and then their performance was re-evaluated with the same test dataset. Table IV shows notable accuracy improvements for all tasks after fine-tuning. For ALC, accuracy increased from 20.71%–50.91% to 86.61%–99.51%. For TSR, accuracy saw a substantial rise from 1.19%–31.24% to 90.34%–99.14%. In the VD task, accuracy improved from 60.64%–92.06% to 88.01%–99.95
@@
These findings suggest that although they initially underperformed on AV tasks, fine-tuning them with relevant datasets can lead to substantial performance improvements, showing their potential utility in AV applications. Intersection over Union (IoU) measures the overlap between the predicted output and the ground truth in tasks like object detection. A higher IoU indicates more accurate localization, making it a crucial metric for evaluating model performance in AV perception tasks. After fine-tuning, the IoU values improved significantly, with LLaVA-7B increasing from 0.041 to 0.964, LLaVA-13B-LoRA from 0.063 to 0.987, MoE-LLaVA from 0.038 to 0.946, MobileVLM from 0.013 to 0.926, Qwen-VL from 0.018 to 0.9578, and NVILA-8B from 0.183 to 0.9768, as shown in Figure 5.
IV-C RQ2: V2LMs Demonstrate Robustness under Attacks
@@
Adversarial attacks pose a serious threat to deep learning based AV perception systems. Prior works have shown that DNN models suffer significant performance degradation across core AV tasks, including TSR, ALC, and VD, as summarized in Table V. For example, the GTSRB CNN model accuracy collapses to just 1.77% under the Shadow Attack [56], and even adversarial training, one of the most prominent defense methods, only modestly improves accuracy to 25.57%. Similarly, for ALC, OpenPilot ALC performance degrades dramatically to 2.50% under the DRP Attack [37], with established defense strategies such as JPEG compression [17], Gaussian noise addition [49], and autoencoder based denoising [16] achieving negligible improvements (around 3%). In VD, YOLOv3 experiences a drastic accuracy drop to 2.52% under RAUCA Attack [57], with no effective defense method proposed.
@@
Building on these findings, we conduct our own robustness evaluation. To provide a fair and rigorous comparison, we evaluated traditional task-specific DNNs—YOLOv5-cls and ViT-small (TSR), CLRerNet (ALC), and YOLOv5-dt (VD)—on benign and previously unseen adversarial datasets (Table III). Our results confirm severe performance degradation, with accuracy reductions of 73.08% and 39.06% for TSR, 40.40% for ALC, and 32.91% for VD, clearly highlighting their vulnerability to novel attacks. In contrast, our evaluations of V2LMs under the same unseen adversarial conditions reveal inherently superior robustness. Specifically, V2LMs experienced substantially smaller accuracy drops: only 8.62%–15.81% for TSR, 4.94%–14.32% for ALC, and 5.23%–7.06% for VD, consistently maintaining high adversarial accuracy without additional defenses.
IV-D RQ3: Tandem V2LMs Provide Similar Performance at Lower Memory Footprint
@@
We evaluate whether one tandem V2LM provides robustness across AV tasks comparable to separate solo V2LMs. To assess this, VLMs are fine-tuned for three tasks, TSR, ALC, and VD, simultaneously using the same prompts. After fine-tuning, both solo and tandem V2LMs were then evaluated on each task to determine its performance on the same test data. Table IV presents the evaluation results for the tandem design in non-adversarial scenarios, where a single V2LM was trained to handle all three tasks simultaneously. The results reveal that the tandem design achieves high accuracy across perception tasks, often matching or surpassing the performance of solo models. This shows the effectiveness of the tandem design in maintaining robust performance across diverse tasks. Table IV further demonstrates the resilience of V2LMs in AV tasks under adversarial conditions. MobileVLM, MoE-LLaVA, LLaVA-7B, NVILA-8B, Qwen-VL, and LLaVA-13B-LoRA allocated 6.03GB, 11.21GB, 13.56GB, 15.23GB, 16.58GB, and 26.15GB of storage, respectively. These results suggest that a single tandem V2LM can generalize well across multiple tasks, providing robust performance comparable to the solo design, which requires 3x more storage for separate models. The tandem design offers a significant advantage in efficiency by keeping similar performance while requiring much less storage.
TABLE V: DNNs accuracy under attacks and defenses.
@@
ALC OpenPilot- ALC DRP-Attack [37] 2.50% JPEG Compression [17] Bit-Depth Reduction [50] Gaussian Noise [49] Median Blurring [50] $\approx$3% (no effective improvement)